OpenStudy (anonymous):

What is the pattern to 7,21,63,189

OpenStudy (anonymous):

7*3 = 21 21*3 = 63 63*3 = 189

OpenStudy (anonymous):

yes you multiply by 3

OpenStudy (charlotte123):

great job! @dhouser